Technical Staff Blog

Last update on .

Emergency Linux Workstation Reboot

Tstaff will be rebooting several department-managed Linux workstations tomorrow, Wednesday morning at 5am. This is part of an ongoing effort to track down some recent filesystem-related problems.

Workstations to be rebooted include all machines in the Sunlab and some other undergrad spaces, in addition to a few faculty and grad student machines. The complete list ...

Last update on .

CIS has reported that this issue has been resolved. If you are still experiencing issues please contact them at the number below.

 

CIS has reported that a significant amount of users are experiencing an outage of Brown email (Gmail). More info from CIS below:

We are aware of a problem with Google Email - Brown Gmail ...

Last update on .

The department list server stopped delivering mail starting at least as far back as yesterday morning (Thursday the 10th). The problem has been resolved and the list server is chewing through the backlog of messages. If you continue to have problems sending email to department lists or receiving listserv messages, email problem.

Last update on .

CIS will be performing a planned switch upgrade on the 3rd floor of the CIT on Thursday, 1/10 from 6:00 AM to 8:00 AM. During this time users on the 3rd floor may experience brief interruptions (less than 5 minutes) of wired, wireless network connectivity and potentially phone usage while the connections ...

Last update on .

CIS will be performing a planned switch upgrade on the 5th floor of the CIT on Tuesday, 1/8 from 6:00 AM to 9:00 AM. During this time users on the 4th and 5th floors may experience brief interruptions (less than 5 minutes) of wired, wireless network connectivity and potentially phone usage while ...

Last update on .

CIS will be performing planned network maintenance to the wireless access points at the CIT and other locations on Wednesday, 12/5, from 6:00 AM to 8:00 AM.  

During this time, the wireless network, (Brown, RLAB, Brown-Guest and Eduroam) will experience an outage lasting less than 5 minutes, at the following locations:

135 ...

Last update on .

Many services were affected by a GPFS filesystem issue this morning.  An NFS server in our GPFS cluster failed, affecting filesystem access generally, and services that rely on filesystem access.  Systems were affected by at least 10am, and possibly much earlier.  The filesystem was restored to normal operation before 2pm.  A VMWare host server also ...

Last update on .

Reports came in yesterday and today of messages being delayed for many hours.  Some were specifically because of account name changes, which was resolved early in the day, yesterday.  While others were delayed due to an after hours service outage, which has also been resolved.

Email services have been returned to normal and all delayed ...

Yearly archives